Opened 7 years ago

Last modified 5 years ago

#374 reopened flight processing

EUFAR10/01 (AIMWETLAB), flight day 238d/2010, Lake Balaton Peninsula

Reported by: iopa Owned by: knpa
Priority: alpha 4 high Milestone:
Component: Archiving Keywords:
Cc: Other processors: emca

Description (last modified by iopa)

Data location: ~arsf/workspace/EUFAR10_01-2010_238d_Lake_Balaton_Peninsula

Data arrived from ARSF via SATA disk Q on 15/09/2010.
However missing IPAS and LiDAR data on disk Q and on thelma delayed unpacking until 06/10/2010.

Scientific objective: Mapping of species composition and stand health of shore vegetation, mapping of shore elevation and inundation structure.

This flight will also be used for EUFAR10-07.
Scientific Objective: High accuracy land use/land cover mapping, forest type mapping and invasive shrub species location

Priority: a4h

PI: A. Zlinszky

This is the second set of data of the split 238c (see http://arsf-dan.nerc.ac.uk/trac/ticket/373 for details)

Note: Data is currently symlinked from ~arsf/arsf_data/2010/flight_data/archive_overflow and when this is resolved the symlinks will need to be re-directed.

Sensors:

  • Eagle (28/03/2011)
  • Hawk (28/03/2011)
  • Leica LIDAR (04/03/2011)
  • Digital Photography (04/03/2011)

Change History (28)

comment:1 Changed 7 years ago by iopa

Applanix and IPAS are common with 238c and need only be processed once.

comment:2 Changed 7 years ago by knpa

  • Description modified (diff)

Added scientific objective.

comment:3 Changed 7 years ago by knpa

  • Description modified (diff)

This flight will also be used for EUFAR10-07.

comment:4 Changed 7 years ago by emca

Processed the camera imagery from raw to tiff

comment:5 Changed 7 years ago by anch

Taking LiDAR and camera

comment:6 Changed 7 years ago by anch

IPAS for 238c covers this flight and has already been processed so I am using this to process LIDAR and tag photos.

comment:7 Changed 7 years ago by anch

Re-processing camera so that we can re tag with corrected event file.

comment:8 Changed 7 years ago by anch

Lidar submitted for delivery check

comment:9 Changed 7 years ago by emca

LiDAR delivery check complete.

Data is good, just a couple of fixes/typos:
Change julian day to 238d on logsheet
PI name incorrectly spelt on logsheet
Readme: dem resolution should be 2.0
Screenshot names - rename to standard name
DEM name - change to reflect new convention (-lidar-ASTER)

comment:10 Changed 7 years ago by anch

Camera ready for checking at:

/users/rsg/arsf/workspace/EUFAR10_01-2010_238d_Lake_Balaton_Peninsula/delivery/20110303/EUFAR10_01

comment:11 Changed 7 years ago by iopa

Checking camera.

comment:12 Changed 7 years ago by iopa

Camera check;

data and delivery are good and ready to deliver.

comment:13 Changed 7 years ago by anch

  • Description modified (diff)

LiDAR and camera sent to PI

comment:14 Changed 7 years ago by iopa

For Hyperspectral processor ; line 24 for Eagle & Hawk should be sliced around halfway as it looks like (when used for other testing) the lens was covered before the sensors were turned off.

comment:15 Changed 7 years ago by emca

  • Other processors set to emca

Starting Hyperspectral processing

comment:16 Changed 7 years ago by anch

  • Summary changed from EUFAR10/01, flight day 238d/2010, Lake Balaton Peninsula to EUFAR10/01 (AIMWETLAB), flight day 238d/2010, Lake Balaton Peninsula

comment:17 Changed 7 years ago by emca

Hyperspectral data is ready for delivery check at
/users/rsg/arsf/workspace/EUFAR10_01-2010_238d_Lake_Balaton_Peninsula/delivery

SCT values:

Line Eagle Hawk
15 +0.03 -0.01
16 -0.02 0.00
17 0.00 +0.02
18 0.00 0.00
19 +0.03 0.00
20 +0.04 +0.02
21 0.00 0.00
22 0.00 +0.02
23 +0.01 0.00
24 0.00 0.00

Notes:
Line 20 contained a bend so both eagle and hawk lines were processed using -l 250 6000 in azspec
As noted above the end of line 24 was removed. Eagle was processed using -l 0 6590 and hawk using -l 0 6200 in azspec.

comment:18 Changed 7 years ago by iopa

Hyperspectral check complete;

nothing to report, ready to go.

comment:19 Changed 7 years ago by iopa

  • Description modified (diff)

Hyperspectral has been dispatched (28/03/2011)

comment:20 Changed 7 years ago by knpa

  • Component changed from Processing: general to Archiving
  • Owner set to knpa
  • Status changed from new to assigned

Project has been r-synced, beginning archiving.

comment:21 Changed 7 years ago by knpa

AIMMS and GRIMM data present

comment:22 Changed 7 years ago by knpa

Tarball created, made available to NEODC (17/05/11)

comment:23 Changed 6 years ago by knpa

Backed-up by NEODC.

Moved to archive disk.

comment:24 Changed 5 years ago by knpa

  • Resolution set to fixed
  • Status changed from assigned to closed

Removed workspace version. Closing ticket.

comment:25 Changed 5 years ago by emca

  • Resolution fixed deleted
  • Status changed from closed to reopened

Hyperspectral reprocessing completed by adbe

Creating delivery as he is on leave this week

comment:26 Changed 5 years ago by emca

Delivery created, ready for checking.

comment:27 Changed 5 years ago by besm

Delivery checked for reprocessing. No problems found.

comment:28 Changed 5 years ago by anhi

Reprocessing hyperspectral delivered to Andras Zlinszky on 12/09/12 on disk 131.

Note: See TracTickets for help on using tickets.